OpenStudy (anonymous):

Solve this system of equations: 2x - y = 7 5x + y = 21 A. (4,1) B. (5,3) C. (3,6) D. (2,-3)

hartnn (hartnn):

Add those 2 equations, what do u get ?

OpenStudy (anonymous):

x=3.5 + 0.5y x = 4.2 + -0.2y @hartnn

hartnn (hartnn):

is that a new question ?

OpenStudy (anonymous):

no thats what i got

OpenStudy (anonymous):

i think the answer is A

hartnn (hartnn):

2x - y = 7 5x + y = 21 adding them 2x-y + 5x +y = 7+21 notice that y cancels out. can you find x now ?

OpenStudy (anonymous):

x=4 @hartnn

hartnn (hartnn):

yes!

hartnn (hartnn):

2(4) - y = 7 find y :)